In the laboratory you dissolve 15.1 g of zinc acetate in a volumetric flask and add water to a total volume of 250 . mL.

What is the molarity of the solution?

_______M
1
Expert's answer
2019-03-12T08:56:29-0400

M(Zn(CH3COO)2) = 183.48 g/mol

n(Zn(CH3COO)2) = m / M = 15.1 / 183.48 = 0.082

c(Zn(CH3COO)2) = n / V = 0.082 / 0.25 = 0.328 M

Answer: c(Zn(CH3COO)2) = 0.328 M